If the back EMF induced in a coil when current changes from 1 ampere to 0 in 1 millisecond, is 5 volts, the self inductance of the coil is:-



(A) 5H



(B) 1H



(C) 5×10^-3 H



(D) 5×10^3 H

1
Expert's answer
2021-11-15T14:03:35-0500

The current changes from 1A to 0A.

We know that ;

"e=-L\frac {dP} {dt}"

e=Voltage =e.m.f

L=Inductance

d=change in current


"\frac {dP} {dt} =[\frac {0-1}{\frac {1} {10^2}}] (-L) =5"


"=\frac {L} {10^{-3}}=5"


"=5\u00d710 ^{-3}H"